OpenStudy (anonymous):

? Solve for x. -3x - 8 + 6x = 7

OpenStudy (anonymous):

First you have to combine like terms so -3x and 6x

OpenStudy (anonymous):

x=15

OpenStudy (anonymous):

Your not supposed to just give the answers im reporting you

OpenStudy (anonymous):

-3x+6x=7+8 3x=15 x=15/3 x=5

OpenStudy (anonymous):

dont be jealous :p

OpenStudy (anonymous):

Im not jealous but your not supposed to just give out the answers because then no one learns anything and thats against the code

OpenStudy (anonymous):

here every one do the same excluding yo.

OpenStudy (anonymous):

A. x = - 7 /5 B. x = - 1 /3 C. x = 5 D. x = 45

OpenStudy (anonymous):

I havent given the answer at all

OpenStudy (anonymous):

it's c x=5

OpenStudy (anonymous):

-3x+6x=7+8 3x=15 x=15/3 x=5